five years ago, my grandfather was five times as old as I was. Three years from now, my grandfather will be three times as old as I will be. how old am I now?
Answer by lwsshak3(11628)

five years ago, my grandfather was five times as old as I was. Three years from now, my grandfather will be three times as old as I will be. how old am I now?
***
5 yrs ago:
x=Grandchild's age
5x=Grandfather's age
..
3 yrs from now:
x+8=Grandchild's age
5x+8=Grandfather's age
..
5x+8=3(x+8)
5x+8=3x+24
2x=16
x=8
..
now:
Grandchild's age=x+5=13
Grandfather's age=5x+5=45